### 20 Feb 2024 v1

# UMBCLU at SemEval-2024 Task 1A and 1C: Semantic Textual Relatedness with and without machine translation

Abstract
This paper describes the system we developed for SemEval-2024 Task 1, “Semantic Textual Relatedness for African and Asian Languages.” The aim of the task is to build a model that can identify semantic textual relatedness (STR) between two sentences of a target language belonging to a collection of African and Asian languages. We participated in Subtasks A and C and explored supervised and cross-lingual training leveraging large language models (LLMs). Pre-trained large language models have been extensively used for machine translation and semantic similarity. Using a combination of machine translation and sentence embedding LLMs, we developed a unified STR model, TranSem, for subtask A and fine-tuned the T5 family of models on the STR data, FineSem, for use in subtask C. Our model results for 7 languages in subtask A were better than the official baseline for 3 languages and on par with the baseline for the remaining 4 languages. Our model results for the 12 languages in subtask C resulted in $1^{st}$ place for Africaans, $2^{nd}$ place for Indonesian, and $3^{rd}$ place for English with low performance for the remaining 9 languages.
UMBCLU at SemEval-2024 Task 1A and 1C: Semantic Textual Relatedness with and without machine translation
Shubhashis Roy Dipta 22footnotemark: 2and Sai Vallurupalli 22footnotemark: 2 {sroydip1,kolli}@umbc.edu Department of Computer Science and Electrical Engineering University of Maryland, Baltimore County Baltimore, MD 21250 USA
footnotetext: These authors contributed equally to this work
1 Introduction
The objective of the SemEval 2024 Task 1 is to build and evaluate models capable of identifying relatedness between sentence pairs from a language. Both sentences in a sentence pair are from a single language taken from a selection of 14 languages belonging to 5 language groups. The task is divided into 3 sub-tasks, where each sub-task is aimed at studying a different type of model training: supervised (Subtask A), unsupervised (Subtask B), and cross-lingual (Subtask C). The languages and the data made available for each subtask are different. Extensive details about the languages, language groups, data, and the collection process can be found in the task description paper Ousidhoum et al. (2024a). A detailed description of the task, sub-tasks, and datasets can be found in the shared task description Ousidhoum et al. (2024b).
Semantic relatedness helps with understanding language meaning (Jarmasz and Szpakowicz, 2012; Miller, 1992) and is useful in many areas of natural language processing such as word-sense disambiguation (Banerjee and Pedersen, 2003), machine translation (Bracken et al., 2017), sentence representation (Reimers et al., 2019), question answering and summarization (Wang et al., 2022) and memory association (Antony et al., 2022; Osgood, 1949). However, semantic relatedness has been mostly restricted to finding word relatedness (Feng et al., 2017; Budanitsky and Hirst, 2006), leading to a lack of sentence relatedness datasets. At the sentence level, relatedness has been limited to similarity, leading to a set of semantic textual similarity (STS) benchmark datasets and models. Similarity is a subset of relatedness and provides a very restricted view of STR (Abdalla et al., 2021). The current shared task aims to develop models and resources for sentence-level relatedness in several languages (Ousidhoum et al., 2024b).
In this paper, we describe our models for the supervised and cross-lingual training subtasks A & C. For subtask A, our aim was to develop a single unified model trained on all the training data for all languages and find the STR of a sentence pair from any language. In such a model, we also could find STR for pairs where each sentence is from a different language. To train the unified model, we use machine translation to convert sentences from all the languages into English. Our target language is English because of the available sentence embedding models in English. For subtask C, our goal was to fine-tune a pre-trained LLM on the STS benchmark dataset and then the English STR dataset and evaluate its usefulness in the cross-domain setting. T5 was a natural choice as the STS dataset is part of the model’s pretraining data.
Our primary contributions in this paper are as follows:
  1. 1.
    We explore if a single unified model can be trained based on translations of the source language.
  2. 2.
    We explore the use of translation models for augmenting data.
  3. 3.
    We explore whether a T5 trained on the English STR dataset works well for any of the target languages in the cross-domain setting.
Our code is publicly available at https://github.com/dipta007/SemEval24-Task8
2 System Overview
After exploring models and datasets available in languages we understand [^1] Of the 14 languages, the authors are proficient in English, Hindi & Telugu languages., we realized the dearth of resources in these languages. Our goal was to explore how to leverage resources available in English for the languages in the shared task. The result of our effort is two different systems for the supervised and cross-lingual training sub-tracks A & C. Our goal for subtask A was to develop a single model for all languages and explore the effectiveness of machine translation. Our goal for subtask C was to explore the effectiveness of transfer learning from English STS to other languages.
2.1 Model for Subtask A
In this section, we describe our translation-based sentence embedding system to use in the supervised setting. Our goal is to translate the sentences from the various languages to English. We train a single model on the translated data and perform inference on the translated data. Assuming the translated data reflects the semantic meaning of the source language accurately, the derived relatedness value from our model reflects the STR between a sentence pair in the source language. The performance of our model allows us to evaluate the machine translation performance by comparing it with the performance of the English dataset, which does not undergo translation.
2.1.1 Translation to English & Data Augmentation
While a number of SOTA translation models are available, we selected Meta’s “No Language Left Behind (NLLB) open-source models” that provide direct high-quality translations for 200 languages where many are low resource languages(Costa-jussà et al., 2022). We used four of the translation models222 facebook/nllb-200-3.3B, facebook/nllb-200-1.3B, facebook/nllb-200-distilled-1.3B, facebook/nllb-200-distilled-600M ranging from 600 million to 3.3 billion parameters for translation to English. We translated the training data for all languages in Subtask A, except Amharic and Algerian Arabic, with each of the models. Thus, we obtained 4 translated datasets for each language, leading to a 4-fold augmentation of our data. We excluded Amharic and Algerian Arabic from our study as none of these models supported the two languages. Aiming to obtain the best translation features at inference time, for the test data, we translated using only the largest model (facebook/all-200-3.3 B).
2.1.2 Our Model (TranSem)
Figure 1: Overview of our TranSem model architecture (Inspired by Reimers et al. (2019)). The encoder ($\theta$) is shared, and the diamond box represents the loss function. The input to the cosine similarity loss is the encoded sentence pairs ($x_{1},x_{2}$) and the label ($y$).
Inspired by Reimers et al. (2019), we used the Siamese model architecture (shown in Fig. 1). For a given pair of sentences ($x_{1},x_{2}$) and their semantic relatedness score ($y$), we encode each sentence with a sentence encoder ($\theta$). The embeddings for the sentences go through a pooling operation ($P$) to produce sentence embeddings ($s_{1},s_{2}$). The cosine similarity of the encoded embeddings is trained to match the semantic relatedness score using the mean-squared error loss:
$\mathcal{L}=MSE(\text{cos-sim}(P\theta(x_{1}),P\theta(x_{2})),y)$ (1)
We experimented with several sentence encoding models and chose DistilRoberta333sentence-transformers/all-distilroberta-v1 to submit results for the competition leaderboard. This is a distilled version of RoBERTa (Liu et al., 2019) fine-tuned on sentence-level datasets and suitable for clustering and semantic searches, which we further fine-tuned on our translated and augmented dataset. We selected this model for submitting the results based on our primary validation (details on § 4.1.1, § 4.1.2, and § 4.1.3). After experimenting with different pooling mechanisms of mean, max, and CLS tokens, we found that mean pooling works well for our setting. This aligns with earlier findings, which showed that mean pooling produced encodings that lead to better performance on downstream tasks.
2.2 Our Model for Subtask C (FineSem)
In this section, we describe our use of the T5 family of models for cross-lingual training. T5 is a text-to-text transformer model that uses transfer learning where the model is pre-trained on a mixture of unsupervised and supervised tasks in a multi-task setting Raffel et al. (2020). One of the supervised tasks used in the T5 model training was the semantic textual similarity benchmark (STS-B) dataset trained as a regression classification problem. We fine-tuned the pre-trained T5-base, T5-large, and XL models ranging from (660 million to 3 billion parameters) on the STR English training data, treating using the same format as the STS-B dataset and using regression classification. Since this model is not allowed for use on the English test set for subtask C (according to the rules for the subtask), we fine-tuned another pretrained-T5 model on the Spanish STR training data for use on the English test set. We also used these fine-tuned T5 models for Subtask A and compared them with our TranSem model.
3 Experimental Setup
3.1 HyperParameters
To train our model, we have used AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 1e-5 and weight decay of 0.01. We use an effective batch size of 32 (batch size 16 with 2 steps of gradient accumulation) (for TransSem) and a batch size of 16 (for FineSem). We train our model infinitely with an early stopping patience of 10 on the validation Spearman Correlation score.
3.2 Infrastructure
All experiments are conducted on an NVIDIA Quadro RTX 8000 with 48GB of VRAM and A100 80GB. We utilize the PyTorch Lightning library  [^4] https://lightning.ai/ for conducting the experiments and Weight & Biases  [^5] https://wandb.ai/ for logging purposes (for the TransSem Model) and HuggingFace Transformers (for the FineSem Model).
4 Results
In this section, first, we report our results and analysis on Subtask A (§ 4.1), and then we report the same on Subtask C (§ 4.2).
4.1 Subtask A
We use the official baseline (described on Ousidhoum et al. (2024b)). They have used LaBSE (Feng et al., 2020) fine-tuned on the provided training dataset. Throughout the paper, we refer to this model as $baseline$.
We have fine-tuned our model (shown in Fig. 1) on the training dataset. Throughout the paper, we refer to this model as $TranSem$.
4.1.1 Effect of Batch Size
We tried various batch sizes (results in Table 1) and showed that our batch size selections are fairly good (32 for task A and 8 for SubTask C).
Batch
Size
A3
eng
A4
hau
A5
kin
A6
mar
A7
ary
A8
esp
A9
tel
avg
TranSem+ 2 .8102 .6857 .6886 .8515 .7376 .6519 .8342 .7514
TranSem+ 4 .5415 .1355 .0643 .5394 .3038 .5998 .4231 .3725
TranSem+ 8 .8132 .6519 .6642 .8348 .7340 .6355 .8187 .7360
TranSem+ 16 .8093 .6377 .6997 .8429 .7500 .6462 .8257 .7445
TranSem+ 64 .8092 .6589 .6456 .8271 .7247 .6234 .8131 .7289
TranSem+ 128 .8129 .6787 .6659 .8417 .7411 .6396 .8324 .7446
TranSem+ 256 .8152 .6716 .6589 .8357 .7197 .6353 .8189 .7365
Table 1: The effect of batch size on $TranSem$ for different batch sizes $\{2,4,8,16,64,128,256\}$
4.1.2 Effect of Encoder Pooling
For the official results we submitted, we used mean pooling as it showed good performance in most languages. In Table 2, we compare results with 3 different pooling operations.
Pool
ing
A3
eng
A4
hau
A5
kin
A6
mar
A7
ary
A8
esp
A9
tel
avg
TranSem+ CLS .8133 .6737 .6655 .8339 .7376 .6363 .8309 .7416
TranSem Mean .8125 .6403 .6807 .8406 .7448 .7211 .8255 .7522
TranSem+ Max .7960 .6157 .5809 .8227 .6643 .6075 .7997 .6981
Table 2: The effect of pooling on $TranSem$ using different pooling mechanisms (CLS Token, Mean, Max)
4.1.3 Effect of Different Models
We provide constrastive results with several other models in Table 3. For the sentence embedding models, we reported official results using distilroberta-v1. We also submitted some comparative results with a T5-XL model fine-tuned on the STR training datasets without any translation (last row in the Table 2)
Model
A3
eng
A4
hau
A5
kin
A6
mar
A7
ary
A8
esp
A9
tel
avg
baseline .8300 .6900 .7200 .8800 .7700 .7000 .8200 .7729
distilroberta-v1 (TranSem) .8125 .6403 .6807 .8406 .7448 .7211 .8255 .7522
mpnet-base-v2 .8104 .6692 .6971 .8568 .7297 .6518 .8250 .7486
roberta-large-v1 .8260 .6750 .7056 .8461 .7480 .6298 .8394 .7528
sentence-t5-xl .8236 .6440 .6720 .8324 .7124 .6277 .8250 .7339
multi-qa-mpnet-base-dot-v1 .8111 .6852 .7041 .8482 .7163 .6586 .8245 .7497
all-MiniLM-L12-v2 .8237 .6474 .7026 .8522 .7605 .6141 .8247 .7465
fine-tuned T5-XL .8381 .6056 .6713 .1827 .7714 .6089 .6818 .6228
Table 3: Model Performance (Spearman Correlation Coefficient) on Subtask A test set. $TranSem$ shows results submitted before the official deadline, $baseline$ shows official baseline results, and the rest are contrastive results for various models we tried. The bold value represents the best score within a specific section, and the underlined value signifies the best score across all sections. The last row “fine-tuned T5-XL” shows the results of a T5-XL model fine-tuned on the STR training data for each language
4.2 Subtask C
We use the English fine-tuned models for inference on all language test sets except English and use the Spanish fine-tuned models for inference on the English test set. We submitted official results using the FineSem model, a T5-XL model fine-tuned for 3 epochs. Our results in Table 4 show the performance of fine-tuned T5-base, Large, and XL models fine-tuned for 10 epochs. We report the best performance for each language, along with the checkpoint saved at a particular epoch number that gave the best performance.
C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11
Model Afr Arq Amh Eng Hau Hin Ind Kin Arb Ary Punj
Baseline .6200 .4600 .8400 .8000 .6200 .7600 .4700 .5700 .6100 .4000 -.050
FineSem .8223 1263 .0430 .7875 .4569 .1552 .5153 .4836 .0354 -1.038 -.078
T5-B .7926 .1799 .2037 .7448 .4784 .2472 .4517 .4437 .0307 .0307 -.308
T5-L .8068 -.035 .0965 .7794 .5052 .2089 .4846 .5529 .0539 .0331 -.012
Table 4: Model performance (Spearman Correlation Coefficient) on Subtask C test sets. All language test sets (except English) use the FineSem models trained on the English training set. The English test set uses the FineSem models trained on the Spanish training set. T5-B, T5-L point to fine-tuning other model sizes of T5 for reporting constrastive results.
5 Related Work
In this section, we will present the previous research conducted in the fields of Machine Translation (see § 5.1), Sentence Embedding (see § 5.2) and Semantic Similarity (see § 5.3).
5.1 Machine Translation
Machine translation has evolved in the last 75 years from rule-based systems to statistical-based systems to the current neural machine translation (NMT) systems. In the 10 years since the first sequence-to-sequence NMT model (Bahdanau et al., 2014), machine translation reached a point where translations from models for high-resource languages rival human translators (Läubli et al., 2018; Popel et al., 2020). This was possible due to the amount of bilingual data pairs available for training in these languages (Haddow et al., 2022). Translation systems for medium and low resource languages that lacked the scale of these resources either developed cross-lingual models (Nguyen and Chiang, 2017; Zoph et al., 2016) or developed datasets (Bañón et al., 2020; Schwenk et al., 2019). Current state-of-the-art translation models use a many-to-many approach to handle a large number of medium to low-resource languages (Costa-jussà et al., 2022).
5.2 Sentence Embedding
Generating a sentence-level embedding is useful for semantic searches and clustering. Since the first transformer model (Vaswani et al., 2017), several encoder-only models such as BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), XLNet (Yang et al., 2018) were used to learn effective sentence embeddings that also performed well on downstream NLP tasks (Cer et al., 2018; Roy Dipta et al., 2023). Reimers et al. (2019) developed a siamese-like network architecture with two BERT sentence embedding models that improved semantic search systems. Using T5 (Raffel et al., 2020) models in a similar architecture, sentence embeddings produced by T5 were shown to be superior to the encoder-only model embeddings with performance gains in downstream tasks Ni et al. (2021). A more recent work (Reimers and Gurevych, 2020) showed that a teacher-student model can be efficiently used to develop a sentence embedding system for many low-resource languages.
5.3 Semantic Similarity
Semantic textual similarity captures a type of semantic relatedness requiring similarity on all aspects between a sentence pair. SemEval tasks on semantic textual similarity from 2012 to 2017 resulted in the STS benchmark (Cer et al., 2017). Recently, Deshpande et al. (2023) proposed conditional semantic textual similarity to explore semantic relatedness.
6 Conclusion & Future Work
We developed two different models and showed how the models performed in supervised and cross-domain training tasks in 14 languages. We explored the use of machine translation, sentence encoders, and SST-B style training with T5 models. Our models improved over the official baseline for some of the languages. For computational purposes, we have excluded using more recent models like mistral-7b (Jiang et al., 2023), which have outperformed most of the open-source and close-source models in various benchmarks (Zheng et al., 2024). Also, we believe that future work can use the prompt-based LLMs  [^6] https://chat.openai.com/ for translation.
7 Limitations
We acknowledge our work has the following limitations. We use several pre-trained large language models in our experiments. It is well known that these models can echo biases and misinformation either implicitly or explicitly. We did not control for any of these when training them on the STR datasets. In addition, the STR datasets may also echo several biases related to social groups, cultural groups, race, gender, behavioral, and perceptual differences of annotators. We did not explore or control for any of these biases in our work. As a result, our work carries the limitations of both the models and the datasets we used.