### 12 Apr 2024 v2

# UMBCLU at SemEval-2024 Task 1: Semantic Textual Relatedness with and without machine translation

Abstract
The aim of SemEval-2024 Task 1, “Semantic Textual Relatedness for African and Asian Languages” is to develop models for identifying semantic textual relatedness (STR) between two sentences using multiple languages (14 African and Asian languages) and settings (supervised, unsupervised, and cross-lingual). Large language models (LLMs) have shown impressive performance on several natural language understanding tasks such as multilingual machine translation (MMT), semantic similarity (STS), and encoding sentence embeddings. Using a combination of LLMs that perform well on these tasks, we developed two STR models, TranSem and FineSem, for the supervised and cross-lingual settings. We explore the effectiveness of several training methods and the usefulness of machine translation. We find that direct fine-tuning on the task is comparable to using sentence embeddings and translating to English leads to better performance for some languages. In the supervised setting, our model performance is better than the official baseline for 3 languages with the remaining 4 performing on par. In the cross-lingual setting, our model performance is better than the baseline for 3 languages (leading to $1^{st}$ place for Africaans and $2^{nd}$ place for Indonesian), is on par for 2 languages and performs poorly on the remaining 7 languages.
UMBCLU at SemEval-2024 Task 1: Semantic Textual Relatedness with and without machine translation
Shubhashis Roy Dipta 22footnotemark: 2and Sai Vallurupalli 22footnotemark: 2 {sroydip1,kolli}@umbc.edu Department of Computer Science and Electrical Engineering University of Maryland, Baltimore County Baltimore, MD 21250 USA
footnotetext: These authors contributed equally to this work
1 Introduction
The objective of the SemEval 2024 Task 1 is to build and evaluate models capable of identifying relatedness between a sentence pair. Sentence pairs from 14 African and Asian languages belonging to 5 language groups are annotated for relatedness and released for model development. The task is divided into 3 tracks targeting different types of model training: supervised (Track A), unsupervised (Track B), and cross-lingual (Track C). Each track targets a different subset of languages. Extensive details about the languages, language groups, and the data collection process are provided in the task description paper Ousidhoum et al. (2024a). A detailed description of the shared task, tracks, and datasets are provided in the shared task description paper Ousidhoum et al. (2024b).
Semantic relatedness helps with understanding language meaning (Jarmasz and Szpakowicz, 2012; Miller, 1995; Antony et al., 2022; Osgood, 1949) and is useful in many areas of natural language processing such as word-sense disambiguation (Banerjee and Pedersen, 2003), machine translation (Bracken et al., 2017) and sentence representation (Reimers et al., 2019; Wang et al., 2022) which have numerous applications. Until recently, semantic relatedness has been mostly restricted to finding word relatedness (Feng et al., 2017; Budanitsky and Hirst, 2006), leading to a lack of sentence-relatedness datasets. At the sentence level, relatedness has been limited to similarity, providing a restricted view of STR (Abdalla et al., 2021). The current shared task aims to broaden the scope of sentence relatedness and extend it to several languages with the goal of encouraging model and resource development in these languages (Ousidhoum et al., 2024b).
Recent advancements in multi-lingual translation and the availability of models for obtaining high-quality sentence embeddings allowed us to explore the effectiveness of machine translated data. Using various sentence embedding models to encode data translated into English, we trained a model, TranSem, to find the relatedness score between sentence pairs. Although the task requires the sentence pair to be from the same language, our model can handle sentences from two different languages. Our second model, FineSem directly fine-tunes a T5 model (T5 is already fine-tuned on the STS benchmark) on the STR task using both untranslated and translated data to explore the usefulness of translation. We use both these models to evaluate languages in Track A. For Track C languages, we use a T5 model fine-tuned only on the english STR data. For evaluating the English dataset in the cross-lingual track, we use a T5 model fine-tuned on the Spanish dataset.
Our contributions to the STR task are as follows: We 1) develop unified models for STR to work with all languages. 2) participate in supervised and cross-lingual tracks. 3) explore the usefulness of machine translation. 4) explore data augmentation using machine translation. Our code is publicly available at https://github.com/dipta007/SemEval24-Task8
2 System Overview
After exploring models and datasets available in the languages we understand [^1] Of the 14 languages, the authors are proficient in English, Hindi & Telugu languages., we realized the dearth of resources available in these languages. To leverage resources available in English, we translated the 13 non-English languages into English. Assuming the translated data accurately reflects the semantic meaning of the source language, the derived relatedness value from our model for a translated sentence pair should reflect the STR between the sentence pair in the source language. This section describes our machine translation process and models, TranSem and FineSem. Besides using different training strategies, these models can use both translated and untranslated data.
2.1 Translation to English & Data Augmentation
We use Meta’s “No Language Left Behind (NLLB) open-source models” that provide direct high-quality translations for 200 languages with many low-resource languages (Costa-jussà et al., 2022). We use four of the translation models222 facebook/nllb-200-3.3B, facebook/nllb-200-1.3B, facebook/nllb-200-distilled-1.3B, facebook/nllb-200-distilled-600M ranging from 600 million to 3.3 billion parameters. Using each of the 4 models, we translated the training data for all languages in track A, except Amharic and Algerian Arabic, and obtained 4 translated datasets for each language. None of the 4 models we used supported Amharic, Algerian Arabic or Punjabi. We decided against translating track C languages with 3 of 12 unsupported languages. Using 4 different model translations gave us a 4-fold augmentation of the training data. We translated the test data using only the largest model (facebook/all-200-3.3 B) to obtain the best translation features.
2.2 TranSem Model
Figure 1: Overview of TranSem model architecture (Inspired by Reimers et al. (2019)). The encoder ($\theta$) is shared, and the diamond box represents the loss function. The encoded sentence pairs ($x_{1},x_{2}$) and the label ($y$) are the input to the cosine similarity loss.
Inspired by Reimers et al. (2019), we used the Siamese model architecture (shown in Fig. 1). For a given pair of sentences ($x_{1},x_{2}$) and their semantic relatedness score ($y$), we encode each sentence with a sentence encoder ($\theta$). The embeddings for the sentences go through a pooling operation ($P$) to produce sentence embeddings ($s_{1},s_{2}$). The cosine similarity of the encoded embeddings is trained to match the semantic relatedness score using the mean-squared error loss:
$\mathcal{L}=MSE(\text{cos-sim}(P\theta(x_{1}),P\theta(x_{2})),y)$ (1)
We experimented with several sentence encoding models for encoding our translated and augmented training dataset. We chose DistilRoberta333sentence-transformers/all-distilroberta-v1 to submit results for the competition leaderboard based on our primary validation (details on § 4.1.1, § 4.1.2, and § 4.1.3). This is a distilled version of RoBERTa (Liu et al., 2019) fine-tuned on sentence-level datasets and suitable for clustering and semantic searches, which we further fine-tuned on our translated and augmented dataset. The sentence-t5-xl embedding model was chosen to compare the effectiveness of sentence embeddings as opposed to the direct fine-tuning used in the FineSem model. After experimenting with different pooling mechanisms of mean, max, and CLS tokens, we found that mean pooling works well for our setting. This aligns with earlier findings, which show that mean pooling produced encodings lead to better performance on downstream tasks.
2.3 FineSem Model
T5 Raffel et al. (2020) is a transformer model that uses transfer learning; the model trained on “Colossal Clean Crawled Corpus” is fine-tuned on a mixture of 8 downstream unsupervised and supervised tasks converting them into a unified text-to-text task setting. The T5 model is available in several sizes, of which we use the base, large, and XL models ranging from (660 million to 3 billion parameters). One of the supervised tasks used in the T5 model training is the semantic textual similarity benchmark (STS-B) dataset trained as a regression classification problem. We use the STS task setting to train on the track A STR training datasets using 3 different options: separate T5 models trained on individual languages, a single model trained on all 14 languages (without translation), and a single model trained on the translated and augmented dataset (for 12 languages). These settings allow us to contrast the effectiveness of direct fine-tuning with the sentence embedding-based transem model and the usefulness of machine translation.
From the T5 models fine-tuned on the individual languages, we use the English and Spanish fine-tuned models (we refer to these models as the English and Spanish models) for evaluating the cross-lingual Track C languages. We use the English model to evaluate development and test data from all languages except English and the Spanish model to evaluate the English data.
3 Experimental Setup
3.1 HyperParameters
We train our models using AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 1e-5 and weight decay of 0.01. We use an effective batch size of 32 (batch size 16 with 2 steps of gradient accumulation) (for TransSem) and a batch size of 16 (for FineSem).
We train our transem model infinitely with an early stopping patience of 10 on the validation Spearman Correlation score. We train the finesem model for 10 epochs (2 epochs for the model trained on the translated and augmented model) and checkpoint the models at the end of every epoch. We evaluate the dev sets for each language against these 10 checkpoints. We evaluate the corresponding test data using the checkpoint which provides the best performance on the dev data for a language.
3.2 Infrastructure
All experiments were conducted on an NVIDIA Quadro RTX 8000 with 48GB of VRAM and A100 80GB. We utilize the PyTorch Lightning library  [^4] https://lightning.ai/ for conducting the experiments and Weight & Biases  [^5] https://wandb.ai/ for logging purposes (for the TransSem Model) and HuggingFace Transformers (for the FineSem Model).
4 Results
We first report our results and analysis on Track A languages (§ 4.1), and then on Track C (§ 4.2). We use the official baseline (Ousidhoum et al. (2024b)) that used LaBSE (Feng et al., 2020) fine-tuned on the provided training dataset and refer to this model as $baseline$.
4.1 Track A Languages
This section discusses our findings using various model settings with the TranSem model.
4.1.1 Effect of Batch Size
Comparing performance with various batch sizes (results are shown in Table 1), we show that our batch size selections are fairly good (32 for task TranSem and 8 for FineSem).
Batch
Size
A3
eng
A4
hau
A5
kin
A6
mar
A7
ary
A8
esp
A9
tel
avg
TranSem+ 2 .8102 .6857 .6886 .8515 .7376 .6519 .8342 .7514
TranSem+ 4 .5415 .1355 .0643 .5394 .3038 .5998 .4231 .3725
TranSem+ 8 .8132 .6519 .6642 .8348 .7340 .6355 .8187 .7360
TranSem+ 16 .8093 .6377 .6997 .8429 .7500 .6462 .8257 .7445
TranSem+ 64 .8092 .6589 .6456 .8271 .7247 .6234 .8131 .7289
TranSem+ 128 .8129 .6787 .6659 .8417 .7411 .6396 .8324 .7446
TranSem+ 256 .8152 .6716 .6589 .8357 .7197 .6353 .8189 .7365
Table 1: The effect of batch size on $TranSem$ for different batch sizes $\{2,4,8,16,64,128,256\}$
4.1.2 Effect of Encoder Pooling
In Table 2, we compare performance using 3 different pooling operations. We used mean pooling for the official results we submitted, as it showed good performance in most languages.
Pool
ing
A3
eng
A4
hau
A5
kin
A6
mar
A7
ary
A8
esp
A9
tel
avg
TranSem+ CLS .8133 .6737 .6655 .8339 .7376 .6363 .8309 .7416
TranSem Mean .8125 .6403 .6807 .8406 .7448 .7211 .8255 .7522
TranSem+ Max .7960 .6157 .5809 .8227 .6643 .6075 .7997 .6981
Table 2: The effect of pooling on $TranSem$ using different pooling mechanisms (CLS Token, Mean, Max)
4.1.3 Effect of Sentence Embedding Models
In Table 3, we provide contrastive results with several sentence embedding models used in TranSem. For official results, we submitted results from the distilroberta-v1 sentence embedding model (results for some of the languages are from the FineSem model fine-tuned on individual STR training datasets).
Model
A3
eng
A4
hau
A5
kin
A6
mar
A7
ary
A8
esp
A9
tel
avg
baseline .8300 .6900 .7200 .8800 .7700 .7000 .8200 .7729
distilroberta-v1 (TranSem) .8125 .6403 .6807 .8406 .7448 .7211 .8255 .7522
mpnet-base-v2 .8104 .6692 .6971 .8568 .7297 .6518 .8250 .7486
roberta-large-v1 .8260 .6750 .7056 .8461 .7480 .6298 .8394 .7528
sentence-t5-xl .8236 .6440 .6720 .8324 .7124 .6277 .8250 .7339
multi-qa-mpnet-base-dot-v1 .8111 .6852 .7041 .8482 .7163 .6586 .8245 .7497
all-MiniLM-L12-v2 .8237 .6474 .7026 .8522 .7605 .6141 .8247 .7465
FineSem-Individual .8385 .6335 .7175 .2211 .7647 .6900 .6085 .6391
FineSem-Unified .8438 .6369 .6837 .3878 .6265 .7040 .6993 .6546
FineSem-Translated .8105 .6383 .7133 .8608 .7403 .6663 .8152 .7493
Table 3: Model Performance (Spearman Correlation Coefficient) on Subtask A test set. $TranSem$ shows results submitted before the official deadline, $baseline$ shows official baseline results, and the rest are contrastive results for our various models. The best scores within each section are bolded, and best scores across all sections are underlined.
4.1.4 Usefulness of Machine Translation and Direct Fine-tuning
We compare the performance of the FineSem models fine-tuned using the 3 data options (results are shown in Table 3). FineSem-Individual shows the performance of T5-XL models fine-tuned on the individual datasets. Unified and Translated models show the performance of the two T5-XL models fine-tuned on the untranslated data and translated$+$augmented data. The model trained on untranslated data performs poorly on the Marathi dataset, but performs on par with the other models indicating that we may not need to translate all languages to English. We find that direct fine-tuning with the translated and augmented data is comparable with the TranSem model using various sentence embeddings.
4.2 Track C Languages
In Table 4, we compare the performance of various T5 models on the track C languages. We submitted official results using our T5-XL based FineSem model (FineSem-LB) where the results are obtained using the checkpoint after the third epoch. With the same model we also report with the approach where we use the checkpoint which results in the best performance on the development data for a given language. These results are shown as FineSem-XL. We compare these results with the T5-base and T5-large based FineSem models. We bold the best scores for easy readability but underline scores that are better than the baseline. Among our models the overall performance of the XL model is better and this model improves upon the baseline for Afrikaans, Indonesian and Spanish.
C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11 C12
Model Afr Arq Amh Eng Hau Hin Ind Kin Arb Ary Pan Esp avg
Baseline .7900 .4600 .8400 .8000 .6200 .7600 .4700 .5700 .6100 .4000 -.050 .6200 .5742
FineSem-LB .8223 1263 .0430 .7875 .4569 .1552 .5153 .4836 .0354 -.0375 -.0775 .6089 .3266
FineSem-XL .8164 1023 .0373 .7889 .4561 .1594 .5279 .4128 .$0$0219 -.0817 .6259 .3246
FineSem-L .8007 -.0515 .0112 .7752 .4831 .1764 .4419 .5094 .0154 .0331 -.0591 .6605 .3164
FineSem-B .7802 .1799 .2543 .7448 .4784 .2404 .4517 .3861 .0527 .0268 -.0520 .6289 .3477
Table 4: Model performance (Spearman Correlation Coefficient) on Track C test sets. All language test sets (except English) use the FineSem models trained on the English training set. The English test set uses the FineSem models trained on the Spanish training set. The best scores among our models are bolded. Scores better than baseline are Underlined.
5 Related Work
In this section, we present previous research conducted in the fields of Machine Translation (see § 5.1), Sentence Embedding (see § 5.2) and Semantic Similarity (see § 5.3).
5.1 Machine Translation
Machine translation has evolved in the last 75 years from rule-based systems to statistical-based systems to the current neural machine translation (NMT) systems. In the 10 years since the first sequence-to-sequence NMT model (Bahdanau et al., 2014), machine translation reached a point where translations from models for high-resource languages rival human translators (Läubli et al., 2018; Popel et al., 2020). This was possible due to the amount of bilingual data pairs available for training in these languages (Haddow et al., 2022). Translation systems for medium and low resource languages that lacked the scale of these resources either developed cross-lingual models (Nguyen and Chiang, 2017; Zoph et al., 2016) or developed datasets (Bañón et al., 2020; Schwenk et al., 2019). Current state-of-the-art translation models use a many-to-many approach to handle a large number of medium to low-resource languages (Costa-jussà et al., 2022).
5.2 Sentence Embedding
Generating a sentence-level embedding is useful for semantic searches and clustering. Since the first transformer model (Vaswani et al., 2017), several encoder-only models such as BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), XLNet (Yang et al., 2018) were used to learn effective sentence embeddings that also performed well on downstream NLP tasks (Cer et al., 2018; Roy Dipta et al., 2023). Reimers et al. (2019) developed a Siamese-like network architecture with two BERT sentence embedding models that improved semantic search systems. Using T5 (Raffel et al., 2020) models in a similar architecture, sentence embeddings produced by T5 were shown to be superior to the encoder-only model embeddings with performance gains in downstream tasks Ni et al. (2021). A more recent work (Reimers and Gurevych, 2020) showed that a teacher-student model can be efficiently used to develop a sentence embedding system for many low-resource languages.
5.3 Semantic Similarity
Semantic textual similarity captures a type of semantic relatedness requiring similarity on all aspects between a sentence pair. SemEval tasks on semantic textual similarity from 2012 to 2017 resulted in the STS benchmark (Cer et al., 2017). Recently, Deshpande et al. (2023) proposed conditional semantic textual similarity to explore semantic relatedness.
6 Conclusion & Future Work
We developed two different models and showed how the models performed in supervised and cross-domain training tasks in 14 languages. We explored using machine translation, sentence encoders, and SST-B style training with T5 models. Our models improved over the official baseline for some of the languages. For computational purposes, we have excluded using more recent models like mistral-7b (Jiang et al., 2023), which have outperformed most of the open-source and close-source models in various benchmarks (Zheng et al., 2024). For future work, we intend to explore prompting for STR and prompt-based LLMs  [^6] https://chat.openai.com/ for translation.
7 Disclaimer
We did not use AI assistants to write any part of our paper or code. All writing is original and produced by the authors.
8 Limitations
We acknowledge our work has the following limitations. We use several pre-trained LLMs in our experiments. It is well known that these models can echo biases and misinformation either implicitly or explicitly. We did not control for any of these when training them on the STR datasets. In addition, the STR datasets may also echo several biases related to social groups, cultural groups, race, gender, behavioral, and perceptual differences of annotators. We did not explore or control for any of these biases in our work. As a result, our work carries the limitations of both the models and the datasets we used.